OpenStudy (anonymous):

7 (6-2a) = 5 (-3a+1)

OpenStudy (anonymous):

7 (6-2a) = 5 (-3a+1) 42 -14a=-15a+5 15a-14a= 4-42 a= -38

OpenStudy (anonymous):

so far I got 42-2a = -15a +1

OpenStudy (anonymous):

@dpasingh where did you get 4 from?

OpenStudy (anonymous):

shouldn't it be 5-42?

OpenStudy (anonymous):

a=-37?

OpenStudy (anonymous):

 Got it ?:)

OpenStudy (anonymous):

yes but schouldn't it be 5-42= -37 not 4-42

OpenStudy (anonymous):

What ???

OpenStudy (anonymous):

7 (6-2a) = 5 (-3a+1) 42 -14a=-15a+5 15a-14a= 4-42 a= -38